OpenStudy (aglna):

The legs of a right triangle are 10 centimeters and 24 centimeters long. What is the length of the hypotenuse? 22 centimeters 26 centimeters 30 centimeters 34 centimeters

OpenStudy (calculusxy):

Use the Pythagorean Theorem: \(a^2 + b^2 = c^2\)

OpenStudy (calculusxy):

a and b are the leg measurements and c is the hypotenuse

OpenStudy (aglna):

I got 26 is that correct

OpenStudy (aglna):

Sorry I meant 34

OpenStudy (calculusxy):

\(10^2 + 24^2 = c^2\) \(676 = c^2\) \(\sqrt{676} = c\) \(\text{Answer: }\bf{26}\)

OpenStudy (calculusxy):

Your initial answer was correct

OpenStudy (aglna):

Ok thanks

OpenStudy (calculusxy):

You're welcome